The 27A current feeding a radio transmitter tower runs straight up the 37m tall tower in an area where the earth magnetic field has a magnitude of 63µT and is pointing downwards at an angle of 23 degrees below due north. What is the magnitude and direction of the force on current? Give direction relative to Up-Down, North-South and/or East-West
Answer #1

F = L*(IxB)
F = L*I*B* sin thetha

angle between I and B is 90 + 23 = 113 degree
To find the direction use Fleming's right hand rule
Index-->current
middle finger-->magnetic field
then thumb gives the direction of force
which will be toward west

F = L*I*B* sin thetha
= 37*27*(63*10^-6)* sin (113)
=0.058 N

Force = 0.058 N toward west